# Teaching Spell Checkers to Teach: Pedagogical Program Synthesis for Interactive Learning

### v1

@@ 5.1. Spell Checker
The spell checker is the entry point of SPIRE, responsible for detecting misspellings and identifying their most likely target words. It operates continuously in the background and is activated after every second long pause in writing. Upon activation, it analyzes the surrounding sentence to locate potential spelling errors and predict their intended forms. For example, when a learner writes “Reechable,” the spell checker identifies the error and generates “Reachable” as the target word. We designed this module to address the limitations of conventional spell checkers, which rely primarily on string similarity and often fail when a learner’s attempt deviates substantially from the correct form (e.g., yous → use). Even commercial tools like Grammarly (grammarly2025) tend to substitute alternative word choices rather than identifying the learner’s intended word. Because SPIRE aims to reason about the learner’s linguistic understanding, it requires target predictions that are both contextually grounded and semantically faithful. To achieve this, the spell checker uses a large language model (LLM) to perform contextual error detection and correction. The LLM considers both the local sentence and the broader discourse context to infer the intended word rather than relying solely on edit distance. Once the target word is identified, both the attempted and target forms—along with their sentence context—are passed to the next module, the word property synthesizer, which begins the process of linguistic analysis.
@@ 5.4.2. Trace Generator & Trace Selector
The trace generator and trace selector together form the core reasoning mechanism of the spelling inquiry engine. Their role is to construct and refine instructional sequences—called traces—that guide the learner through 2–5 hypothesis-driven steps of exploration. The generator proposes multiple alternative instructional paths, while the selector evaluates and ranks these paths to identify the one most pedagogically sound. This two-phase design is modeled after human inductive reasoning and draws inspiration from the hypothesis search framework (wang2023hypothesis), which separates divergent generation from convergent evaluation. Effective instruction requires both creativity and precision: educators first imagine several possible ways to guide a learner, then choose the one that best advances understanding. By mirroring this cognitive process, SPIRE balances exploration (diversity of ideas) with instructional rigor (quality of reasoning).
@@ 5.5. Interface Generator
We designed this module to preserve an isomorphic relationship between pedagogical logic and interface behavior. Rather than displaying static feedback, the interface visually manifests the structure of the instructional program—showing learners not just the correction, but the reasoning process behind it. This design enables the same program to operate seamlessly in both headless (text-based) and interactive (graphical) modes, making it adaptable across different learning contexts. Technically, the interface generator parses the JSON execution plan produced by the program synthesizer and compiles it into React components. Each component corresponds to a hypothesis in the instructional trace (e.g., h1, h2, h3), specifying its instructional text, input affordances (such as text fields or visual tools like paint-rollers), and verification conditions. Conditional branches (on(true), on(false)) define how the interface adapts to learner responses in real time, ensuring pedagogical continuity across different interaction outcomes. This direct coupling between program logic and user interaction enables SPIRE to render the reasoning process of an SLP as an interactive, scaffolded learning experience.
@@ 5.6. Implementation Details
SPIRE integrates a Python backend with a React frontend. The spelling inquiry engine is implemented in Python using a unification-based pattern matcher (maclellan_pyplan) and Anthropic’s Claude Sonnet (anthropic_claude_sonnet45) for hypothesis filtering, generation, and selection. The synthesized instructional programs are serialized as JSON execution plans, which the React interface renders as interactive components mirroring the system’s pedagogical logic. To support multimodality, SPIRE employs Google Gemini’s speech-to-text for capturing spoken responses and Gemini’s text-to-speech for verbalizing instructions and feedback, enabling voice-based interaction alongside text input. This modular setup ensures consistent execution across conversational and interactive modes while remaining adaptable for integration with future writing or tutoring systems.
@@ 6.1. Material
We collected 10 writing samples from children in grades 2-5, each containing one or more spelling errors. Using the SPIRE pipeline, we generated 25 conversations between a simulated expert tutor and simulated student, with each conversation addressing one misspelled word. Samples with multiple errors were corrected across successive conversations. Each conversation included 2-5 just-in-time interventions. For each intervention, the system provided: (1) the targeted SWI question type (meaning, structure, relatives, or grapheme-phoneme correspondence), (2) the selected hypothesis for testing, (3) the rationale for this selection, and (4) the instructional dialogue. The system’s responses dynamically adjusted to simulated learner responses following the branching logic expressed in our domain-specific language.
@@ 6.2. Participants
We recruited 5 experts through word-of-mouth and mailing lists. The demographic information pertinent to each expert are present in the table 3. The number of participants is in line with prior work that has required domain-experts/teachers (gong2025callisense; salminen2024deus).
6.3. Study Protocol
Experts first attended a Zoom orientation session explaining the evaluation task. They then rated conversation transcripts through a custom web interface using 5-point Likert scales for each intervention: (Q1) “The action selected is appropriate”; (Q2) “The reason for action selection is valid”; (Q3) “I feel confident in my assessment.” Experts could also add qualitative notes. After completing one practice transcript with guided walkthrough, participants independently rated all 25 transcripts over multiple days (2-6 hours total). Participants received \$50 compensation.
@@ 6.4. Results
6.4.1. Pedagogical Reasoning Validity.
Experts demonstrated exceptionally high consensus on the soundness of the system’s pedagogical reasoning (Q1), with 78.9% perfect agreement across all five raters ($\mu=4.93$, $\sigma=0.41$, average item $\sigma=0.14$). The near-ceiling ratings indicate strong expert agreement that the system consistently provided appropriate pedagogical explanations for student errors.
@@
Actions received consistently positive ratings ($\mu=4.38$, $\sigma=1.15$) with moderate inter-rater agreement (23.9% perfect agreement). The higher variance (average item $\sigma=0.78$) reflects the pedagogically complex nature of tutoring, where multiple valid instructional approaches often exist for addressing student errors. Qualitative notes revealed that disagreement primarily centered on: (1) pedagogical sequencing (e.g., whether to address morphology before grapheme-phoneme correspondence), (2) depth of explanation appropriate for different age groups, and (3) terminology preferences (e.g., “base” vs. “base word,” “vowel digraph” vs. “vowel team”).
6.4.3. Expert Confidence.
Raters reported high confidence in their evaluations ($\mu=4.93$, $\sigma=0.41$, 78.9% perfect agreement), validating the reliability of the expert assessment process.
@@
The contrasting patterns of expert agreement reveal two important characteristics of the system. The high agreement on reasoning validity (78.9% perfect agreement in Table 4) demonstrates that the system exhibits strong pedagogical competence—it consistently generates sound rationales that experts recognize as appropriate. In contrast, the moderate agreement on instructional actions (23.9% perfect agreement in Table 4) reflects genuine pedagogical flexibility rather than a system limitation. Effective tutoring inherently admits multiple valid approaches, and expert teachers frequently disagree on optimal instructional sequences even while concurring on underlying pedagogical principles.
[…]
E1 71 20+ years 8 years
E2 34 9 years 2 years
E3 55 15 years 4 years
E4 47 15 years 8 years
E5 47 20+ years 7 years
@@
[…]
Reasoning 4.93 (0.41) 98.7% 95.4% t=109.93* 78.9%
Action 4.38 (1.15) 83.1% 70.5% t=28.14* 23.9%
Expert 4.93 (0.41) 98.7% 95.4% t=109.93* 78.9%
Table 4. Inter-rater agreement and rating distributions across evaluation dimensions. N = 545 ratings (109 responses × 5 experts per dimension). “% $\geq$ 4” shows the proportion of ratings at 4 (agree) or 5 (strongly agree); “% = 5” shows the proportion at 5 only. t-tests assessed whether mean ratings exceeded the neutral midpoint (3 = “neither agree nor disagree”). *p < .001.
@@ 7.2. Study Procedure
We conducted 45-minute remote sessions via Zoom with children and their parents present. After obtaining parental consent and child assent, we began with a brief demonstration of the tool’s features. Children then selected a word to investigate from a set of five words (selected from the from High Frequency Words Project (loveless2024high) )—friend, definitely, favorite, because, beautiful—to explore using our spelling tool. During the intervention, children were provided with a pre-written sample where their selected word was intentionally misspelled for their morphological investigation.
@@ 7.3.1. Engagement and Perceived Value
Children expressed enthusiasm for continued use of the tool (median = 4, range = 3–5). More significantly, during the retrospective think-aloud, all seven participants spontaneously discussed how the tool could help them learn words beyond just fixing spelling errors. A fourth grade participant shared, “I can learn new words with this!”, referring to the word-matrix mini-game he had to complete while learning about “because” and how its base “cause” can be used to make many new words like “causal” and “causes”. Parents echoed this enthusiasm, with one parent observing, “At school I don’t think he gets taught this way.”
@@ 7.3.4. Interface Design
Children appreciated how the tool guided them through instructions autonomously, making the experience feel more dynamic than filling out a worksheet. Observing the incorrect spelling instance transform into the correct spelling helped them in most instances understand why their hypothesis regarding the spelling was wrong. They liked the different modalities available for input—draggable icons, audio input, and text fields. However, a parent suggested having more animations in the interface to make it more playful and fun for children.
These preliminary findings suggest that the tool addresses a genuine gap in current spelling instruction and demonstrates baseline usability for the target age group. While our sample size prevents statistical claims, the consistent positive responses and identification of morphological learning as novel indicate the tool’s potential for broader impact.
@@ 8.1. Pedagogical Program Synthesis for Personalized and Individualized Learning
Traditional rule-based systems for instructional support perform well in domains with well-defined constraints and limited solution spaces, such as mathematical problem solving, where errors follow predictable patterns. However, literacy instruction, particularly spelling instruction grounded in morphological, phonological, and etymological analysis, presents fundamentally different challenges. The cognitive processes expert educators employ when diagnosing spelling errors and constructing instructional responses cannot be adequately captured through explicit rules alone. A single misspelling can involve multiple overlapping dimensions: grapheme-phoneme mappings, morphological boundaries, suffixing conventions, semantic appropriateness, or etymological patterns. The space of pedagogically valid responses is not enumerable in advance, and the optimal instructional sequence depends on nuanced inference about the learner’s underlying reasoning. This inference requires flexible interpretation rather than pattern matching against predefined templates.
Pedagogical Program Synthesis addresses this limitation by combining the structured reasoning of domain-specific languages with the interpretive flexibility of large language models. The DSL provides instructional scaffolding, encoding hypothesis templates, evidential features, and compositional primitives that preserve pedagogical validity, while LLMs enable dynamic adaptation to individual learners’ contexts. This architecture opens possibilities that purely corrective systems cannot achieve. A learner initially seeking to correct a single grapheme error may discover, through synthesized inquiry sequences, the morphological family structure explaining that grapheme choice, the etymological origins preserving it across related words, and the systematic patterns governing similar cases. The learning effect extends far beyond the immediate correction, building metalinguistic awareness that may be transferred to future encounters.
Current classroom instruction rarely provides this form of structured, individualized morphological investigation. SPIRE demonstrates that Pedagogical Program Synthesis can serve as a vessel for structured literacy experiences that do not currently exist at scale, capturing the dynamic and responsive processes that expert educators employ while making them accessible within everyday writing contexts. The approach suggests a broader applicability to instructional domains that require flexible and context-sensitive reasoning about learner understanding.
@@ 8.2. Clinical Implications
SPIRE demonstrates broad applicability across instructional settings aligned with Multi-Tiered Systems of Supports frameworks commonly implemented in schools. The system architecture enables deployment at multiple intervention levels: classroom teachers can integrate it during universal instruction (Tier 1), specialists can use it for targeted small group support (Tier 2), and speech-language pathologists can incorporate it within intensive individualized intervention (Tier 3). Beyond educator-mediated instruction, the system supports independent student practice, facilitating distributed practice across various learning contexts, a condition known to enhance skill generalization beyond clinical settings.
@@ 8.3. Limitations and Future Work
Our empirical evaluation, while providing valuable insight into the pedagogical validity and usability of SPIRE, involved modest sample sizes: five speech-language pathology experts and seven children. The expert evaluation yielded strong evidence of pedagogical soundness, with near-ceiling ratings and high inter-rater agreement on reasoning validity. The preliminary user study with children revealed consistent patterns that suggested good usability and engagement. However, these sample sizes exclude robust quantitative claims about effectiveness or generalizability in diverse populations of learners, grade levels, and ranges of spelling proficiency. A larger-scale study would strengthen statistical power and enable a more nuanced analysis of how SPIRE’s effectiveness varies between learner characteristics.

### v2

@@ 5.1. Spell Checker
The spell checker is the entry point of SPIRE, responsible for detecting misspellings and identifying their most likely target words. It operates continuously in the background and is activated after every second long pause in writing. Upon activation, it analyzes the surrounding sentence to locate potential spelling errors and predict their intended forms. For example, when a learner writes “Reechable,” the spell checker identifies the error and generates “Reachable” as the target word. We designed this module to address the limitations of conventional spell checkers, which rely primarily on string similarity and often fail when a learner’s attempt deviates substantially from the correct form (e.g., yous → use). Even commercial tools like Grammarly (Grammarly Inc., 2025) tend to substitute alternative word choices rather than identifying the learner’s intended word. Because SPIRE aims to reason about the learner’s linguistic understanding, it requires target predictions that are both contextually grounded and semantically faithful. To achieve this, the spell checker uses a large language model (LLM) to perform contextual error detection and correction. The LLM considers both the local sentence and the broader discourse context to infer the intended word rather than relying solely on edit distance. Once the target word is identified, both the attempted and target forms—along with their sentence context—are passed to the next module, the word property synthesizer, which begins the process of linguistic analysis.
@@ 5.4.2. Trace Generator & Trace Selector
The trace generator and trace selector together form the core reasoning mechanism of the spelling inquiry engine. Their role is to construct and refine instructional sequences—called traces—that guide the learner through 2–5 hypothesis-driven steps of exploration. The generator proposes multiple alternative instructional paths, while the selector evaluates and ranks these paths to identify the one most pedagogically sound. This two-phase design is modeled after human inductive reasoning and draws inspiration from the hypothesis search framework (Wang et al., 2023), which separates divergent generation from convergent evaluation. Effective instruction requires both creativity and precision: educators first imagine several possible ways to guide a learner, then choose the one that best advances understanding. By mirroring this cognitive process, SPIRE balances exploration (diversity of ideas) with instructional rigor (quality of reasoning).
@@ 5.5. Interface Generator
We designed this module to preserve an isomorphic relationship between pedagogical logic and interface behavior. Rather than displaying static feedback, the interface visually manifests the structure of the instructional program—showing learners not just the correction, but the reasoning process behind it. This design enables the same program to operate seamlessly in both headless (text-based) and interactive (graphical) modes, making it adaptable across different learning contexts. Technically, the interface generator parses the JSON execution plan produced by the program synthesizer and compiles it into React components. Each component corresponds to a hypothesis in the instructional trace (e.g., h1, h2, h3), specifying its instructional text, input affordances (such as text fields or visual tools like paint-rollers), and verification conditions. Conditional branches (on(true), on(false)) define how the interface adapts to learner responses in real time, ensuring pedagogical continuity across different interaction outcomes. This direct coupling between program logic and user interaction enables SPIRE to render the reasoning process of an SLP as an interactive learning experience.
@@ 5.6. Implementation Details
SPIRE integrates a Python backend with a React frontend. The spelling inquiry engine is implemented in Python using a unification-based pattern matcher (MacLellan, 2024) and Anthropic’s Claude Sonnet (Anthropic, 2025) for hypothesis filtering, generation, and selection. From text analysis through intervention generation, the pipeline makes 8 LLM API calls total, including parallel calls for hypothesis search and final DSL synthesis; using Claude Sonnet 4.5, this costs approximately \$0.70 per complete word inquiry. The complete pipeline—from text analysis to intervention generation—involves 8 LLM API calls, including parallel threads for hypothesis search. During testing with Sonnet 4.5, this averaged \$0.70 per complete word inquiry (excluding optional speech services). The synthesized instructional programs are serialized as JSON execution plans, which the React interface renders as interactive components mirroring the system’s pedagogical logic. To support multimodality, SPIRE employs Google Gemini’s speech-to-text for capturing spoken responses and Gemini’s text-to-speech for verbalizing instructions and feedback, enabling voice-based interaction alongside text input. This modular setup ensures consistent execution across conversational and interactive modes while remaining adaptable for integration with future writing or tutoring systems.
@@ 6.1. Material
We collected 20 writing samples from children in grades 2-5, each containing one or more spelling errors. Using the SPIRE pipeline, we generated 50 conversations between a simulated expert tutor and simulated student, each addressing one misspelled word with 2-5 just-in-time interventions. To ensure the evaluation workload remained manageable for the experts, the 50 conversations were randomly divided into two sets of 25 distinct transcripts. For each intervention, the system provided: (1) the targeted SWI question type, (2) the selected hypothesis, (3) the rationale, and (4) the instructional dialogue. The system’s responses dynamically adjusted to simulated learner responses following the branching logic expressed in our domain-specific language.
@@ 6.2. Participants
We recruited 10 experts through word-of-mouth and mailing lists. The demographic information pertinent to each expert are present in the table 3. The number of participants is in line with prior work that has required domain- experts/teachers (Gong et al., 2025; Salminen et al., 2024).
6.3. Study Procedure
Experts first attended a Zoom orientation session explaining the evaluation task. They then rated conversation transcripts through a custom web interface using 5-point Likert scales for each intervention: (Q1) “The action selected is appropriate”; (Q2) “The reason for action selection is valid”; (Q3) “I feel confident in my assessment.” Experts could also add qualitative notes. After completing a guided transcript, participants independently rated 25 transcripts from their randomly assigned set over multiple days. Completing the task required 2-6 hours. Participants received \$50 compensation.
@@ 6.4. Results
Experts responded to 109 interventions in the first set and 105 in the second set. We report results jointly for both sets that were given to 5 participants each. This is because the words present between the two sets were of similar difficulty. Additionally, there were no statistical difference across the sets.
6.4.1. Pedagogical Reasoning Validity (Q1).
Experts demonstrated exceptionally high consensus on the soundness of the system’s pedagogical reasoning, with 74.7% perfect agreement across all ten raters ($\mu=4.90$, $\sigma=0.47$). The near-ceiling ratings indicate strong expert agreement that the system consistently provided appropriate pedagogical explanation for student errors.
@@
Actions received consistently positive rating ($\mu=4.32$, $\sigma=1.13$) with moderate inter-rater agreement of 26.7% perfect agreement. The higher variance reflects the pedagogically complex nature of tutoring, where multiple valid instructional approaches often exist for addressing student errors. Qualitative notes revealed that disagreement primarily centered on three topics: (1) 1) pedagogical sequencing (e.g., whether to address morphology before grapheme-phoneme correspondence), (2) depth of explanation appropriate for different age groups, and (3) terminology preferences (e.g., “base” vs. “base word,” “vowel digraph” vs. “vowel team”).
6.4.3. Expert Confidence (Q3).
Raters reported high confidence in their evaluations ($\mu=4.92$, $\sigma=0.47$), along with high agreement of 80.1%, validating the reliability of the expert assessment process.
@@
The contrasting patterns of expert agreement reveal two important characteristics of the system. The high agreement on reasoning validity (74.7% perfect agreement) in Table 4) demonstrates that the system exhibits strong pedagogical competence—it consistently generates sound rationales that experts recognize as appropriate. In contrast, the moderate agreement on instructional actions reflects genuine pedagogical flexibility rather than a system limitation. Effective tutoring inherently admits multiple valid approaches, and expert teachers frequently disagree on optimal instructional sequences even while concurring on underlying pedagogical principles.Experts also noted that the simulated dialogue mirrored their own teaching strategies, praising the flexibility and breadth of generated interventions. Raters specifically commended the system’s orthographic precision, such as correctly identifying the <-ion> suffix in “action” rather than the common error <-tion>.
ID Age Spelling Instruction SWI Experience Experience E1 71 20+ years 8 years E2 34 9 years 2 years E3 55 15 years 4 years E4 47 15 years 8 years E5 47 20+ years 7 years E6 56 14 years 7 years E7 51 20+ years 10 years E8 34 9 years 2 years E9 56 13 years 5 years E10 60 15 years 13 years
@@
Dimension Mean (SD) % $\geq$ 4 % = 5 t-test vs Neutral ($\mu$=3) Perfect Agreement Reasoning 4.90 (0.47) 97.2% 94.4% 41.97* 74.7% Action 4.32 (1.13) 82.0% 70.2% 9.59* 26.7% Confidence 4.92 (0.47) 97.5% 95.7% 43.04* 80.1%
* p < 0.001
Table 4. Inter-rater agreement and rating distributions across evaluation dimensions. % $\geq$4 and % =5 indicate proportions rating "agree" or higher and "strongly agree," respectively. Participant-level t-tests compare means against the neutral midpoint ($\mu$=3).
@@ 7.2. Study Procedure
We conducted 45-minute remote sessions via Zoom with children and their parent present. After obtaining parental consent and child assent, we began with a brief demonstration of the tool’s features. Children then selected a word to investigate from a set of five words (selected from the from High Frequency Words Project (Loveless and Hamilton, 2024) )—friend, definitely, favorite, because, beautiful—to explore using our spelling tool. During the intervention, children were provided with a pre-written sample where their selected word was intentionally misspelled for their morphological investigation.
@@ 7.3.1. Engagement and Perceived Value
Children expressed enthusiasm for continued use of the tool (median = 4, range = 3–5). More significantly, during the retrospective think-aloud, all seven participants spontaneously discussed how the tool could help them learn words beyond just fixing spelling errors. A fourth grade participant shared, “I can learn new words with this!”, referring to the word-matrix mini-game he had to complete while learning about “because” and how its base “cause” can be used to make many new words like “causal” and “causes”. Parents echoed this enthusiasm, with one parent observing, “At school I don’t think he gets taught this way.” Another parent remarked, “He started with one [misspelling] but walked away with so many more words.” This aligns with a key goal highlighted by experts in our technical evaluation: utilizing structured word inquiry to facilitate the acquisition of multiple related words from a single investigation.
@@ 7.3.4. Interface Design
Children appreciated how the tool guided them through instructions autonomously, making the experience feel more dynamic than filling out a worksheet. Observing the incorrect spelling instance transform gradually into the correct spelling helped them in most instances understand why their hypothesis regarding the spelling was wrong. They liked the different modalities available for input—draggable icons, audio input, and text fields. However, a parent suggested having more animations in the interface to make it more playful and fun for children.
These findings suggest that the tool addresses a gap in current spelling instruction and demonstrates baseline usability for the target age group. While our sample size prevents statistical claims, the consistent positive responses and identification of morphological learning as novel indicate the tool’s potential for broader impact.
@@ 8.1. Pedagogical Program Synthesis for Personalized and Individualized Learning
Traditional rule-based systems for instructional support perform well in domains with well-defined constraints and limited solution spaces, such as mathematical problem solving (Siddiqui et al., 2024), where errors follow predictable patterns. However, literacy instruction, particularly spelling instruction grounded in morphological, phonological, and etymological analysis, presents different challenges. The cognitive processes expert educators employ when diagnosing spelling errors and constructing instructional responses cannot be adequately captured through explicit rules alone (Fournier-Viger et al., 2010). A single misspelling can involve multiple overlapping dimensions: grapheme-phoneme mappings, morphological boundaries, suffixing conventions, semantic appropriateness, or etymological patterns. The space of pedagogically valid responses is not enumerable in advance, and the optimal instructional sequence depends on nuanced inference about the learner’s underlying reasoning. This inference requires flexible interpretation rather than pattern matching against predefined templates.
Pedagogical Program Synthesis addresses this limitation by combining the structured reasoning of domain-specific languages with the interpretive flexibility of large language models. The DSL provides instructional scaffolding, encoding hypothesis templates, evidential features, and compositional primitives that preserve pedagogical validity, while LLMs enable dynamic adaptation to individual learners’ contexts. Critically, the DSL also serves as a guardrail against the hallucination risks inherent in generative systems by constraining interventions to valid pedagogical templates and evidence-backed linguistic features. The architecture bounds the space of possible outputs to coherent responses, even when the underlying model produces imperfect completions. This architecture opens possibilities that purely corrective systems cannot achieve. A learner initially seeking to correct a single grapheme error may discover, through synthesized inquiry sequences, the morphological family structure explaining that grapheme choice, the etymological origins preserving it across related words, and the systematic patterns governing similar cases. The learning effect extends far beyond the immediate correction, building metalinguistic awareness that may be transferred to future encounters. From a deployment perspective, the current implementation relies on commercial LLM APIs, which raises legitimate concerns about cost and latency in school settings. However, these constraints are architectural rather than fundamental. Asynchronous background generation and prompt caching can reduce perceived latency. Additionally, future work could explore WebGL-enabled browser-based models for low-cost deployment.
Current classroom instruction rarely provides this form of structured, individualized morphological investigation. SPIRE demonstrates that Pedagogical Program Synthesis can serve as a vessel for structured literacy experiences that do not currently exist at scale, capturing the dynamic processes that expert educators employ while making them accessible within everyday writing contexts. The approach suggests a broader applicability to instructional domains that require flexible and context-sensitive reasoning about learner understanding.
@@ 8.2. Clinical Implications
SPIRE’s architecture is designed to align with the Multi-Tiered Systems of Supports (California Department of Education, 2024) frameworks commonly implemented in schools. While our current evaluation focused on usability and expert validation, the system’s modular design supports potential deployment at multiple intervention levels: classroom teachers could integrate it during universal instruction (Tier 1), specialists could use it for targeted small group support (Tier 2), and speech-language pathologists could incorporate it within intensive individualized intervention (Tier 3). Beyond educator-mediated instruction, the system supports independent student practice, facilitating distributed practice across various learning contexts, a condition known to enhance skill generalization beyond clinical settings.
@@ 8.3. Limitations and Future Work
Our empirical evaluation, while providing valuable insight into the pedagogical validity and usability of SPIRE, involved modest sample sizes: ten speech-language pathology experts and seven children. The expert evaluation yielded strong evidence of pedagogical soundness, with near-ceiling ratings and high inter-rater agreement on reasoning validity. The preliminary user study with children revealed consistent patterns that suggested good usability and engagement. However, these sample sizes exclude robust quantitative claims about effectiveness or generalizability in diverse populations of learners, grade levels, and ranges of spelling proficiency. In particular, future work should include more students who are not native English speakers and whose native language may have a different grammatical structure, such as Chinese, Russian or Spanish. A larger-scale study would strengthen statistical power and enable a more nuanced analysis of how SPIRE’s effectiveness varies between learner characteristics.